OpenStudy (anonymous):

Solve using the addition principle. -6 = x – 2

OpenStudy (anonymous):

x = -4

OpenStudy (anonymous):

-6 = x - 2 +6 - 6 = x - 2 +6 x - 2 + 6 = 0 x = -4

OpenStudy (precal):

-6=x-2 +2 +2 -4=x